4. What is the critical angle for a ruby-air interface if rubys refraction index is 1.54? (2 marks) Page 3
Answer #1

The critical angle between ruby-air interface = gif.latex?%5CTheta

gif.latex?%5CTheta= Sin-1(n2/n1)

where n1 = refletive index of ruby = 1.54

n2= reflective index of air =1

gif.latex?%5CTheta= Sin-1(1/1.54) = 40.48 degrees